OpenStudy (bgrg007):

someone plz explain this to me! find the exact value of this expression. cot^-1(-sqrt(3)). i know the answer is 5pi/6 but i don't understand how that came to be.

OpenStudy (anonymous):

cot = cos/sin so you need to find the angle that gives -sqrt(3) for cos/sin cos(5pi/6) = -sqrt(3)/2 and sin(5pi/6) = 1/2 =>cos(5pi/6)/sin(5pi/6)= (-sqrt(3)/2)/(1/2)=(-sqrt(3)/2)*(2/1) = -sqrt(3) 11pi/6 also works cos(11pi/6) = sqrt(3)/2m and sin(11pi/6) = -1/2

OpenStudy (bgrg007):

thanks bro! :) really clear explanation. :)

OpenStudy (anonymous):

you're welcome1 glad you like it!